# Geometric integrators for adiabatically closed simple thermodynamic systems

### v1

@@ 5 Conclusions and outlook
Inspired by the geometric framework of almost cosymplectic manifolds derived in [d.B2025], we have developed a discrete variational principle for adiabatically closed simple thermodynamic systems, leading to the discrete thermodynamic Euler–Lagrange equations (21). We have also presented a Noether’s theorem for these systems (Theorem 12) as well as for the continuous counterpart (Theorem 3 and Theorem 5. We have illustrated the effectiveness of our method with three examples: the damped harmonic oscillator and both a perfect gas and a Van der Waals gas confined by a piston contained in a cylinder.
@@ Acknowledgements
M. de León and J. Bajo received financial support from Grant PID2022-137909NB-C21 funded by MICIU/AEI/ 10.13039/501100011033 and Grant CEX2023-001347-S funded by MICIU/AEI/10.13039/501100011033. J. Bajo received the support of a fellowship from “La Caixa” Foundation (ID 100010434) with code LCF/BQ/PFA25/11000023.

### v2

@@ 5 Conclusions and outlook
Inspired by the geometric framework of almost cosymplectic manifolds derived in [d.B2025], we have developed a discrete variational principle for adiabatically closed simple thermodynamic systems, leading to the discrete thermodynamic forced Euler–Lagrange equations (22). We have also presented a Noether’s theorem for these systems (Theorem 13) as well as for the continuous counterpart (Theorem 3 and Theorem 5. We have illustrated the effectiveness of our method with four examples: the damped harmonic oscillator, both a perfect gas and a Van der Waals gas confined by a piston contained in a cylinder, and a Van der Waals gas subject to external forces.
@@ Acknowledgements
The authors wish to express their gratitude to the anonymous referees for their constructive comments. M. de León and J. Bajo received financial support from Grant PID2022-137909NB-C21 funded by MICIU/AEI/ 10.13039/501100011033 and Grant CEX2023-001347-S funded by MICIU/AEI/10.13039/501100011033. J. Bajo received the support of a fellowship from “La Caixa” Foundation (ID 100010434) with code LCF/BQ/PFA25/11000023.
Data availability
The Matlab codes employed for the simulations are available upon reasonable request.
Conflict of Interest
The authors declare no conflict of interest.